# DFDT: Dynamic Fast Decision Tree for IoT Data Stream Mining on Edge Devices

### v1

@@ 3 Methodology
12:         $C3\leftarrow H_{l}\geq\text{avg}(H_{LH_{\text{stat}}})-\sigma(H_{LH_{\text{stat}}})$
13:         $C4\leftarrow H_{l}\geq\text{avg}(H_{\text{stat}})-\sigma(H_{\text{stat}})$
14:         $C5\leftarrow G_{\text{best}}\geq\text{avg}(G_{\text{stat}})-\sigma(G_{\text{stat}})$
15:         $C6\leftarrow n_{l}\geq\text{avg}(n_{\text{stat}})$
16:         Update $H_{stat}$ with $H_{l}$, $G_{stat}$ with $G_{best}$, and $n_{stat}$ with $n_{l}$
17:         if C3 $and$ C4 $and$ C5 $and$ C6 then
18:              return True
19:         end if
20:     end if
21:     return False
22:end procedure
The four primary constraints governing the split operate on the instance count, global entropy, historical entropy, and historical information gain. The instance count constraint ensures that splits only occur when the leaf has accumulated sufficient data, by checking if the current instance count $n_{l}$ at the leaf $l$ meets or exceeds the average instance count recorded across all leaves under VFDT-satisfied conditions:
$n_{l}\geq\text{avg}(n_{\text{stat}})$ (3)
Conversely, the remaining three constraints are evaluated using the same general formula $\varphi(x,X)$, which determines that the current metric value $x$ at a leaf is significantly different from reference values $X$ when bigger than the mean of the values in $X$ minus a standard deviation:
@@
Together, these constraints allow DFDT to restrict tree growth in nodes of moderate activity to instances where statistically significant differences exist. For highly active nodes, two skipping conditions [3] are calculated according to the general formula $\omega(x,X)$, which permits a split when the current metric $x$, i.e. the entropy $H_{l}$ and information gain $G_{best}$, is at least one standard deviation above the average historical values in $X$:
@@
These splitting conditions work as a skipping mechanism to the conservative constraints. In these cases, splits are allowed to proceed directly if the current entropy and information gain are significantly higher than historical averages. Subsequently, the statistics for entropy, information gain, HB, and the number of instances seen are updated (Alg. 2, Step 16). All of these are constant-time operations. If the conditions are satisfied, a split is triggered (Alg. 2, Step 18). This split involves navigating the tree, sorting features, and creating new nodes, with the added complexity of re-initializing feature estimators for each branch (Alg. 1, Step 17-21). The time complexity becomes $O(F\log F+B\cdot F)$, where $B\cdot F$ represents the initialization cost for the new branches. If the split attempt fails at any point (Alg. 2, Step 21), the algorithm resets the instance count of the last checked split (Alg. 1, Step 23), and recalculates the grace period $n_{\text{min}}$ based on specific conditions at each leaf (Alg. 1, Steps 24-28). This decision is based on the tie threshold $\tau$, the difference in information gain between the top attributes $\Delta G$ and the statistical confidence bound $\epsilon$, derived from the HB [8], with two possible scenarios. In the first scenario, $\Delta G$ is smaller than $\epsilon$, but still exceeds $\tau$, suggesting that more data is needed to confirm the observed gain. In this case, $n_{\text{min}}$ is increased to allow additional data accumulation, ensuring that the gain difference becomes statistically significant in the subsequent evaluation. In the second scenario, $\Delta G$ is below $\tau$, indicating that the attributes are statistically similar, but $\epsilon$ is still above $\tau$. Here, $n_{\text{min}}$ is further increased to delay the split until more data is gathered, allowing $\epsilon$ to shrink sufficiently. Thus:
@@
$n_{\text{min}}=\begin{cases}\left\lceil\frac{R^{2}\ln(1/\delta)}{2(\Delta G)^{2}}\right\rceil,&\text{if }\Delta G<\epsilon\text{ and }\Delta G>\tau\\ \\ \left\lceil\frac{R^{2}\ln(1/\delta)}{2\tau^{2}}\right\rceil,&\text{if }\Delta G<\tau\text{ and }\epsilon>\tau\end{cases}$ (6)
@@
where $\delta$ is a user-defined confidence parameter, and $R$ is the range of heuristic values. This dynamic adjustment minimizes unnecessary computations and improves energy efficiency, while maintaining the model’s accuracy. Since split attempts (Alg. 1, Steps 14–30) occur only once every $n_{\text{min}}$ instances, their overall contribution to the time complexity is weighted by the grace period factor. Thus, the total time complexity is: $N\cdot O(\log_{B}|LH|+F)+\frac{N}{n_{\text{min}}}\cdot O(F\log F+B\cdot F)$.
4 Experiments
This section evaluates the performance of the Dynamic Fast Decision Tree (DFDT) against the VFDT, svfdt_I, and svfdt_II, while performing an ablation study on the proposed algorithm. The following notation is used to described the isolated contributions of each individual component of the algorithm: additional splitting conditions ($DFDT$), adaptive expansion nodes ($\text{VFDT}_{E}$), adaptive tie threshold ($\text{VFDT}_{T}$), and adaptive grace period ($\text{VFDT}_{G}$). With $\text{VFDT}_{GT}$, $\text{DFDT}_{GE}$, $\text{DFDT}_{GT}$, $\text{DFDT}_{TE}$, and $\text{DFDT}_{GTE}$ being the respective combinations. All models and algorithm components were implemented using a custom experimental environment built with the pystream library [^1] https://github.com/vturrisi/pystream, a Python-based tool for data stream mining, inspired by the Massive Online Analysis (MOA) framework. It supports both Python and Cython implementations, with the latter significantly reducing runtime overhead by compiling Python code into C.
For evaluation, a diverse set of real-world datasets obtained from the USP Data Stream Repository were used [15]. These are shown in Table 1, encompassing both binary and multiclass classification tasks. The experiments were conducted using a prequential (test-then-train) evaluation strategy [7], monitoring the interrelationships among accuracy, memory usage, and runtime. To provide a more intuitive and holistic assessment, an efficiency metric was computed as the equal-weighted sum of these three components. Each component was normalized on a per-dataset basis to ensure consistency and comparability across datasets of varying scales. To maintain a consistent interpretation, the normalized memory and runtime values were inverted, such that lower values correspond to better performance in these dimensions.
Table 1: Experimental datasets
Name No. of classes No. of features No. of examples
Chess 3 7 534
Keystroke 4 10 1,600
Luxembourg 2 30 1,901
@@ 4 Experiments
NOAA Weather 2 8 18,159
SmartMeter LADPU 10 96 22,950
Electricity 2 8 45,312
Rialto Bridge Timelapse 10 27 82,250
@@
Forest Covertype 7 54 581,012
Poker-hand 10 11 829,201
All algorithms were optimized for each dataset using a grid search strategy, with grace periods set to 100, 400, and 1000, and tie-splitting thresholds configured to 0.01, 0.05, and 0.1 as the hyperparameter settings. Across all tables, datasets are presented in ascending order based on the number of instances. Table 2 shows the $\text{DFDT}_{GTE}$’s efficiency against VFDT, svfdt_I and svfdt_II.
Table 2: Efficiency
Dataset $\text{DFDT}_{GTE}$ VFDT svfdt_I svfdt_II
Chess 0.58 0.67 0.64 0.64
Keystroke 0.9 0.66 0.66 0.66
Luxembourg 0.86 0.89 0.84 0.84
Ozone 0.99 0.86 0.87 0.87
NOAA 0.94 0.75 0.75 0.78
SmartMeter 0.74 0.66 0.66 0.66
Electricity 0.63 0.45 0.52 0.51
Rialto 0.69 0.63 0.58 0.64
Posture 0.77 0.71 0.67 0.7
KDDCUP99 0.57 0.47 0.49 0.48
ForestCoverType 0.67 0.62 0.59 0.61
PokerHand 0.52 0.37 0.42 0.4
Overall, $\text{DFDT}_{GTE}$ emerges as a well-rounded top performer, excelling in accuracy, memory usage, and runtime, with the exception of the Chess and Luxembourg datasets. This performance discrepancy is likely due to the limited number of instances, which hinders the full effectiveness of the adaptive mechanisms. Indeed, across all experiments conducted on the four smaller datasets, each containing fewer than 2,500 instances, no consistent performance trends were observed across the algorithms. Consequently, the subsequent analysis focuses on the remaining eight larger real-world datasets. To provide a more nuanced perspective on the trade-offs between accuracy and memory usage across algorithms, Figure 1 illustrates the memory usage and accuracy metrics for each dataset.
@@
Figure 1: Prequential accuracy and memory usage
As shown, $\text{DFDT}_{GTE}$ maintains consistent high performance, outperforming other algorithms in nearly all datasets except ForestCoverType. Although it does not achieve the highest accuracy in this dataset, the difference in accuracy is minimal, and $\text{DFDT}_{GTE}$ demonstrates significantly better memory efficiency. Across the other datasets, $\text{DFDT}_{GTE}$ generally occupies more memory than competing algorithms, reflecting a trade-off for its enhanced accuracy. This balance between memory usage and accuracy suggests that $\text{DFDT}_{GTE}$ is optimized for scenarios where accuracy is prioritized, yet it remains reasonably efficient in memory-constrained environments. To further study this claim, an ablation study is performed on $\text{DFDT}_{GTE}$’s components, with the obtained accuracy across the largest real-world datasets, is presented in Table 3.
Table 3: Prequential accuracy
Algorithm NOAA SmartMeter Electricity Rialto Posture ForestCoverType PokerHand
VFDT 73.1 $\pm$ 2.2 52.4 $\pm$ 2.3 80.3 $\pm$ 1.5 35.0 $\pm$ 4.9 50.5 $\pm$ 2.1 79.5 $\pm$ 2.5 74.2 $\pm$ 4.5
VFDTG 72.6 $\pm$ 2.0 52.4 $\pm$ 2.1 79.4 $\pm$ 1.6 34.1 $\pm$ 4.3 50.7 $\pm$ 1.5 78.3 $\pm$ 2.4 75.4 $\pm$ 4.5
VFDTGT 75.2 $\pm$ 0.1 60.8 $\pm$ 1.8 82.0 $\pm$ 1.1 61.9 $\pm$ 6.0 58.3 $\pm$ 0.5 85.3 $\pm$ 0.2 92.5 $\pm$ 2.3
VFDTT 75.3 $\pm$ 0.6 59.9 $\pm$ 1.2 82.3 $\pm$ 1.0 62.0 $\pm$ 5.5 58.7 $\pm$ 0.7 88.7 $\pm$ 1.4 93.0 $\pm$ 2.1
DFDTE 73.2 $\pm$ 1.2 52.6 $\pm$ 2.6 79.3 $\pm$ 0.7 34.8 $\pm$ 4.1 49.0 $\pm$ 1.7 70.3 $\pm$ 2.5 69.7 $\pm$ 4.1
DFDTG 71.2 $\pm$ 2.0 52.3 $\pm$ 1.9 79.0 $\pm$ 0.6 32.1 $\pm$ 3.4 48.8 $\pm$ 1.7 74.6 $\pm$ 2.0 69.7 $\pm$ 2.0
DFDTGE 72.4 $\pm$ 1.8 52.3 $\pm$ 1.9 79.1 $\pm$ 0.9 33.9 $\pm$ 4.1 48.4 $\pm$ 1.1 72.7 $\pm$ 1.9 69.0 $\pm$ 1.5
DFDTGT 74.6 $\pm$ 0.5 54.8 $\pm$ 0.4 79.0 $\pm$ 0.4 41.7 $\pm$ 0.6 51.8 $\pm$ 1.0 80.1 $\pm$ 1.1 84.7 $\pm$ 2.7
DFDTGTE 75.1 $\pm$ 0.3 54.8 $\pm$ 0.4 79.4 $\pm$ 0.6 42.3 $\pm$ 0.5 52.3 $\pm$ 0.6 70.9 $\pm$ 1.3 72.9 $\pm$ 5.4
DFDTT 74.9 $\pm$ 0.3 59.3 $\pm$ 1.0 80.8 $\pm$ 1.4 51.2 $\pm$ 7.1 56.1 $\pm$ 0.9 84.4 $\pm$ 2.8 85.9 $\pm$ 4.1
DFDTTE 75.0 $\pm$ 0.4 59.8 $\pm$ 0.6 80.0 $\pm$ 1.3 46.6 $\pm$ 4.1 53.0 $\pm$ 0.4 70.1 $\pm$ 2.6 72.4 $\pm$ 0.9
svfdt_I 72.7 $\pm$ 2.1 52.3 $\pm$ 2.0 79.7 $\pm$ 1.0 33.3 $\pm$ 4.6 49.3 $\pm$ 1.8 75.7 $\pm$ 4.1 72.4 $\pm$ 4.2
svfdt_II 73.3 $\pm$ 2.4 52.3 $\pm$ 2.0 80.5 $\pm$ 1.7 34.3 $\pm$ 5.0 50.2 $\pm$ 2.0 77.5 $\pm$ 3.5 74.8 $\pm$ 4.0
Looking at Table 3, we see however that despite $\text{DFDT}_{GTE}$ having better accuracy than VFDT, svfdt_I and svfdt_II, it seems that exclusively adding adaptive tie threshold $\text{VFDT}_{T}$ or also adaptive grace period $\text{VFDT}_{GT}$ yields better results. This hints that most of the predictive performance improvement comes from these mechanisms, while the additional splitting constraints and adaptive expansion modes favor memory usage. Inspecting the accuracy evolution of the largest dataset PokerHand provides further light on how accuracy and memory usage evolve as more instances are processed. Figure 2 is organized into three sections: an ablation of the three adaptive mechanisms while using the additional splitting constraints, an ablation of the two adaptive parameters while using the original splitting condition, and a direct comparison between svfdt_I, svfdt_II, and DFDT.
(a) Ablation of adaptive mechanisms with splitting constraints
(b) Ablation of adaptive parameters with original HB condition
(c) Comparison of svfdt_I, svfdt_II, and DFDT
Figure 2: Prequential accuracy and memory over time.
The first pair of graphs demonstrates the impact of incorporating the adaptive tie threshold $\text{DFDT}_{T}$. This is particularly evident as more data instances are processed. However, this improvement comes at the cost of increased memory usage growing rapidly. To mitigate this issue, introducing the adaptive grace period $\text{DFDT}_{GT}$ helps control tree growth, effectively reducing memory usage while only slightly affecting accuracy. Further enhancements, such as the adaptive expansion modes with leaf deactivation $\text{DFDT}_{GTE}$, yield even more substantial memory savings, however substantially hindering accuracy. The second pair of graphs shows that adaptive tie threshold has the same positive impact on accuracy and memory when using the original splitting condition without the additional constraints. Finally, the third pair of graphs highlights the advancements brought by the proposed algorithm. While svfdt_I and svfdt_II have a more stable performance, they lack the predictive and resource efficiency of DFDT with the four combined mechanisms: additional splitting conditions, adaptive expansion nodes, adaptive tie threshold, and adaptive grace period. For a big picture of this trade-off between accuracy and memory, their Nemeyi critical difference diagrams are shown in Figures 4 and 4, respectively.
Figure 3: Accuracy
Figure 4: Memory
While VFDT implementations with adaptive parameters consistently achieve high ranks, above 4, in terms of accuracy, this improved performance comes with a trade-off of increased tree growth, adversely impacting memory usage. Conversely, algorithms using the adaptive tie threshold along with the additional splitting constraints, adaptive expansion modes and/or adaptive grade period have significantly better memory usage. To further understand how the different combinations of components affect this trade-off, Table 4 revisits the aggregate efficiency ranking across real-world datasets.
Table 4: Evaluation ranking
Algorithm Efficiency Accuracy Memory Runtime
$\text{DFDT}_{GT}$ 0.78 0.54 0.91 0.93
$\text{DFDT}_{GTE}$ 0.74 0.43 0.94 0.87
$\text{VFDT}_{GT}$ 0.72 0.86 0.47 0.84
$\text{VFDT}_{G}$ 0.7 0.28 0.95 0.89
$\text{DFDT}_{G}$ 0.67 0.18 0.97 0.87
$\text{DFDT}_{GE}$ 0.67 0.19 0.97 0.86
VFDT 0.65 0.29 0.92 0.74
svfdt_II 0.65 0.29 0.93 0.74
svfdt_I 0.64 0.23 0.95 0.75
$\text{DFDT}_{T}$ 0.61 0.56 0.59 0.72
$\text{DFDT}_{TE}$ 0.58 0.38 0.66 0.73
$\text{VFDT}_{T}$ 0.52 0.8 0.0 0.77
$\text{DFDT}_{E}$ 0.44 0.31 0.94 0.08
Overall, the results reveal that the hoeffding tree with adaptive parameters ($\text{VFDT}_{GT}$) excels in scenarios where accuracy is paramount, but may fall short in memory-constrained environments. Thus, careful consideration must be given to the memory overhead introduced, especially in extreme edge environments. On the other hand, the proposed algorithm with all the various adaptive components ($\text{DFDT}_{GTE}$) strikes a compelling balance between accuracy, memory usage and runtime. However, omitting adaptive expansion modes ($\text{DFDT}_{GT}$) seems to lead to even superior overall efficiency, by increasing accuracy and reducing computation time. The choice of algorithm configuration should be guided by the specific demands of the deployment environment. If memory constraints are a significant concern, the full $\text{DFDT}_{GTE}$ is recommended. For scenarios where maximum accuracy is paramount, and memory usage is less of an issue excluding the additional splitting constraints and/or adaptive expansion nodes alone may suffice.
5 Conclusions

### v2

@@ Introduction
12:   $C3\leftarrow H_{l}\geq\overline{H}_{LH_{\text{stat}}}-\sigma(H_{LH_{\text{stat}}})$
13:   $C4\leftarrow H_{l}\geq\overline{H}_{\text{stat}}-\sigma(H_{\text{stat}})$
14:   $C5\leftarrow G_{\text{best}}\geq\overline{G}_{\text{stat}}-\sigma(G_{\text{stat}})$
15:   $C6\leftarrow n_{l}\geq\overline{n}_{\text{stat}}$
16:   Update $H_{stat}$ with $H_{l}$
17:   Update $G_{stat}$ with $G_{best}$
18:   Update $n_{stat}$ with $n_{l}$
19:   if C3 andand C4 andand C5 andand C6 then
20:     return True
21:   end if
22:  end if
23:  return False
24:end procedure
The four primary constraints governing the split operate on the instance count, global entropy, historical entropy, and historical information gain. The instance count constraint ensures that splits only occur when the leaf has accumulated sufficient data, by checking if the current instance count $n_{l}$ at the leaf $l$ meets or exceeds the mean instance count recorded across all leaves under VFDT-satisfied conditions. Conversely, the remaining three constraints are evaluated using the same general formula $\varphi(x,X)$, which determines that the current metric value $x$ at a leaf is significantly different from reference values $X$ when bigger than the mean of the values in $X$ minus a standard deviation:
@@
Together, these constraints allow DFDT to restrict tree growth in nodes of moderate activity to instances with statistically significant differences. For highly active nodes, two skipping conditions are calculated according to $\omega(x,X)$, which permits a split when the current metric $x$, i.e. the entropy $H_{l}$ and information gain $G_{best}$, is at least one standard deviation above the mean historical values in $X$:
@@
These splitting conditions work as a skipping mechanism to the conservative constraints. In these cases, splits are allowed to proceed directly if the current entropy and information gain are significantly higher than historical means. Subsequently, the statistics for entropy, information gain, HB, and the number of instances seen are updated (Alg. 2, Step 16-18). All of these are constant-time operations. If the conditions are satisfied, a split is triggered (Alg. 2, Step 18). This split involves navigating the tree, sorting features, and creating new nodes, with the added complexity of re-initializing feature estimators for each branch (Alg. 1, Step 19-25). The time complexity becomes $O(F\log F+B\cdot F)$, where $B\cdot F$ represents the initialization cost for the new branches. If the split attempt fails at any point (Alg. 2, Step 23), the algorithm resets the instance count of the last checked split (Alg. 1, Step 28), and recalculates the grace period $n_{\text{min}}$ based on specific conditions at each leaf (Alg. 1, Steps 29-32). This decision is based on the tie threshold $\tau$, the difference in information gain between the top attributes $\Delta G$ and the statistical confidence bound $\epsilon$, derived from the HB (garcia2018hoeffding), with two possible scenarios. In the first scenario, $\Delta G$ is smaller than $\epsilon$, but still exceeds $\tau$, suggesting that more data is needed to confirm the observed gain. In this case, $n_{\text{min}}$ is increased to allow additional data accumulation, ensuring that the gain difference becomes statistically significant in the subsequent evaluation. In the second scenario, $\Delta G$ is below $\tau$, indicating that the attributes are statistically similar, but $\epsilon$ is still above $\tau$. Here, $n_{\text{min}}$ is further increased to delay the split until more data is gathered, allowing $\epsilon$ to shrink sufficiently. Thus:
@@
$n_{\text{min}}=\begin{cases}\left\lceil\frac{R^{2}\ln(1/\delta)}{2(\Delta G)^{2}}\right\rceil,&\text{if }\tau<\Delta G<\epsilon\\ \\ \left\lceil\frac{R^{2}\ln(1/\delta)}{2\tau^{2}}\right\rceil,&\text{if }\Delta G<\tau<\epsilon\end{cases}$ (4)
@@
where $\delta$ is a user-defined confidence parameter, and $R$ is the range of heuristic values. This adjustment minimizes unnecessary computations and improves energy efficiency, while maintaining the model’s accuracy. Since split attempts (Alg. 1, Steps 16–35) occur only once every $n_{\text{min}}$ instances, their overall contribution to the time complexity is weighted by the grace period factor. Thus, the total time complexity is: $N\cdot O(\log_{B}|LH|+F)+\frac{N}{n_{\text{min}}}\cdot O(F\log F+B\cdot F)$.
Experiments
All models and algorithm components were implemented within the pystream library [^1] https://github.com/vturrisi/pystream, a Cython-based tool for data stream mining, reducing runtime overhead by compiling Python code into C. For evaluation, a diverse set of real-world datasets obtained from the USP Data Stream Repository [^2] https://sites.google.com/view/uspdsrepository were used (Souza2020). These are shown in Table 2, encompassing both binary and multiclass classification tasks. The experiments were conducted using a prequential (test-then-train) evaluation strategy (gama2009issues), monitoring the interrelationships among accuracy, memory usage, and runtime. All algorithms were optimized for each dataset using a grid search strategy, with grace periods ($n_{\text{min}}$) set to 100, 400, and 1000, and tie-splitting thresholds ($\tau$) configured to 0.01, 0.05, and 0.1 as the hyperparameter settings. The Hoeffding bound confidence level was fixed at $1\times 10^{-7}$ across all experiments. DFDT’s adaptive control of grace periods and tie thresholds removes the need for tuning these specific hyperparameters, as they govern tree growth and thus become implicitly determined by the HB confidence. However, DFDT introduces two fixed activity thresholds. While this raises questions about their adaptivity, it is important to note that, in contrast to $n_{\text{min}}$ and $\tau$, adapting these activity thresholds through the HB confidence is inappropriate, because activity operates conditionally on HB-driven splits and modulates leaf deactivation in ways that counteract tree growth. In practice, consistent with findings from GAHT (garcia2022green), we observed that a leaf-deactivation threshold below 2% of the expected instance count ($f_{\text{deactivate}}=0.02$) and a skip-check threshold at twice the expected count ($f_{\text{expand}}=2$) provide robust performance across datasets. Majority class prediction was used at the leaf level for all models.
Name Classes Features Examples
@@
NOAA 2 8 18,159
METER 10 96 22,950
ELEC 2 8 45,312
RIALTO 10 27 82,250
@@
COVER 7 54 581,012
POKER 10 11 829,201
Table 2: Experimental datasets
Figure 2: Main and interaction effects of DFDT components
Since DFDT integrates multiple heuristics (Rules, Activity, $\tau$, and $n_{\text{min}}$), it is essential to conduct an ablation study to assess the individual and combined contributions of these components to runtime, memory usage, and accuracy. This analysis helps identify Pareto-efficient variants of the algorithm. To ensure fair comparison across datasets of varying scales, all results were normalized on a per-dataset basis. For interpretability, normalized memory and runtime values were inverted, so that higher scores consistently reflect better performance across all metrics. Figure 2 presents coefficient estimates from linear regression models, each fitted separately for the three target outcomes. Interaction terms were included to capture the synergistic effects between heuristics. Individually, the adaptive grace period, activity-aware pruning, and stricter splitting rules did not produce statistically significant accuracy–memory–runtime trade-offs. Moreover, the combinations of Rules × Activity and Rules × $n_{\text{min}}$ also demonstrated stable trade-offs, informing the design of $\text{DFDT}_{\text{Low}}$ and $\text{DFDT}_{\text{Medium}}$, respectively. In contrast, the interaction between Activity × $n_{\text{min}}$ and the inclusion of an adaptive $\tau$ significantly improved accuracy, albeit at the expense of memory efficiency. Additionally, $\tau$’s interactions with other heuristics exhibited inconsistent effects on accuracy, suggesting a more complex dynamic. As a result, $\text{DFDT}_{\text{High}}$ activates all heuristics to maximize accuracy, potentially at the cost of runtime or memory consumption.
Model Rules Activity $n_{\text{min}}$ $\tau$
$\text{DFDT}_{\text{\text{Low}}}$ X X
$\text{DFDT}_{\text{\text{Medium}}}$ X X
$\text{DFDT}_{\text{\text{High}}}$ X X X X
Table 3: Pareto-efficient DFDT variants
Figures 1 and 3 show the average performance across all datasets, plotting prequential accuracy against worst-case memory usage (MB) and per-instance computational time (µs), respectively. Table 4 reports the detailed results for each dataset. Among the baseline methods, VFDT exhibits moderate predictive performance while being the slowest and most memory-intensive model, underscoring its inefficiency across both metrics. Its modified variant, VFDT-$n_{\text{min}}$, offers a modest reduction in memory consumption and a notable improvement in runtime, with minimal impact on accuracy. SVFDT demonstrates the lowest memory usage and maintains competitive runtime with VFDT-$n_{\text{min}}$, however this efficiency is achieved at the expense of predictive accuracy. In contrast, the proposed DFDT variants delineate a clear Pareto frontier with respect to memory usage and share a competitive frontier with VFDT-$n_{\text{min}}$ in terms of runtime. Specifically, $\text{DFDT}_{\text{Low}}$ stands out as the most resource-efficient model, exhibiting lower runtime and memory consumption than SVFDT while maintaining comparable accuracy, making it particularly suitable for deployment in resource-constrained environments. The $\text{DFDT}_{\text{Medium}}$ variant achieves a favorable trade-off, significantly improving accuracy while preserving strong efficiency in both runtime and memory. Finally, $\text{DFDT}_{\text{High}}$ attains the highest overall accuracy, accompanied by a modest increase in computational time and a more pronounced increase in memory usage. Thus, only appropriate when resource constraints are less critical.
Figure 3: Accuracy x Runtime
To assess the statistical significance of the observed differences, we applied the Friedman test followed by the Nemenyi post-hoc test. The resulting Critical Difference Diagrams (CDDs), shown in Figures 4, 5, and 6, display the algorithm rankings at a 95% confidence level. In these diagrams, lower values correspond to better average ranking across datasets. The CDD for accuracy reveals that $\text{DFDT}_{\text{High}}$ achieves the best overall rank, and is the only method that significantly outperforms SVFDT. This result confirms the effectiveness of incorporating all heuristic components within DFDT to maximize predictive performance.
Figure 4: Nemenyi test - Accuracy
Method NOAA METER ELEC RIALTO POSTURE COVER POKER Avg. (Rank)
VFDT
73.1 $\pm$ 2.2
0.47 MB
43.2 µs
52.4 $\pm$ 2.3
14.19 MB
288.5 µs
80.3 $\pm$ 1.5
0.14 MB
49.8 µs
35.0 $\pm$ 4.9
3.37 MB
118.1 µs
50.5 $\pm$ 2.1
0.22 MB
78.5 µs
79.5 $\pm$ 2.5
0.80 MB
254.1 µs
74.2 $\pm$ 4.5
0.43 MB
963.4 µs
63.6 (2.79)
2.80 (5.36)
256.5 (5.57)
[…]
72.6 $\pm$ 2.0
0.31 MB
35.0 µs
52.4 $\pm$ 2.1
10.37 MB
285.1 µs
79.4 $\pm$ 1.6
0.10 MB
38.4 µs
34.1 $\pm$ 4.3
2.87 MB
104.7 µs
50.7 $\pm$ 1.5
0.21 MB
69.3 µs
78.3 $\pm$ 2.4
0.62 MB
241.2 µs
75.4 $\pm$ 4.5
0.43 MB
916.8 µs
63.3 (3.14)
2.13 (4.14)
241.5 (3.00)
SVFDT
71.2 $\pm$ 2.0
0.15 MB
33.0 µs
52.3 $\pm$ 1.9
3.73 MB
283.0 µs
79.0 $\pm$ 0.6
0.05 MB
35.7 µs
32.1 $\pm$ 3.4
0.87 MB
109.0 µs
48.8 $\pm$ 1.7
0.10 MB
69.0 µs
74.6 $\pm$ 2.0
0.39 MB
255.3 µs
69.7 $\pm$ 2.0
0.38 MB
905.3 µs
61.1 (5.29)
0.81 (1.93)
241.5 (2.21)
$\text{DFDT}_{\text{Low}}$
72.4 $\pm$ 1.8
0.16 MB
33.3 µs
52.3 $\pm$ 1.9
2.76 MB
279.1 µs
79.1 $\pm$ 0.9
0.07 MB
37.8 µs
33.9 $\pm$ 4.1
0.54 MB
110.2 µs
48.4 $\pm$ 1.1
0.09 MB
70.3 µs
72.7 $\pm$ 1.9
0.47 MB
239.8 µs
69.0 $\pm$ 1.5
0.37 MB
907.8 µs
61.1 (5.21)
0.64 (1.71)
239.8 (2.64)
$\text{DFDT}_{\text{Medium}}$
75.1 $\pm$ 0.3
0.21 MB
35.3 µs
54.7 $\pm$ 0.5
2.92 MB
254.7 µs
79.4 $\pm$ 0.6
0.10 MB
39.0 µs
42.3 $\pm$ 0.5
0.47 MB
112.3 µs
52.3 $\pm$ 0.6
0.20 MB
69.4 µs
70.9 $\pm$ 1.3
1.12 MB
260.5 µs
72.9 $\pm$ 5.4
1.23 MB
942.8 µs
63.9 (2.43)
0.89 (3.71)
246.4 (4.14)
$\text{DFDT}_{\text{High}}$
74.6 $\pm$ 0.5
0.21 MB
34.0 µs
54.8 $\pm$ 0.4
12.19 MB
280.5 µs
79.0 $\pm$ 0.4
0.05 MB
36.2 µs
41.7 $\pm$ 0.6
3.21 MB
105.7 µs
51.8 $\pm$ 1.0
0.20 MB
70.3 µs
80.1 $\pm$ 1.1
1.05 MB
255.3 µs
84.7 $\pm$ 2.7
1.23 MB
942.9 µs
66.7 (2.14)
2.59 (4.14)
244.9 (3.42)
@@
Table 4: Prequential accuracy (%), memory usage (MB), and per-instance computational time (µs)
The ranking for memory consumption exhibits an inverse trend compared to accuracy. VFDT is the most memory-intensive algorithm, receiving the worst average rank. In contrast, $\text{DFDT}_{\text{Low}}$ achieves the best overall rank. Notably, VFDT performs significantly worse than both $\text{DFDT}_{\text{Low}}$ and SVFDT, whereas $\text{DFDT}_{\text{Medium}}$ and $\text{DFDT}_{\text{High}}$ do not differ significantly, highlighting their well rounded performance.
Figure 5: Nemenyi test - Memory
The CDD for runtime identifies VFDT as the slowest algorithm. Among the compared methods, only SVFDT demonstrated statistically significant better performance than VFDT. Although the variant $\text{DFDT}_{\text{Low}}$ exhibited a lower mean runtime than SVFDT, its ranking was not consistently superior across datasets.
Figure 6: Nemenyi test - Runtime
Conclusions